# Masked Thought: Simply Masking Partial Reasoning Steps Can Improve Mathematical Reasoning Learning of Language Models

### v1

@@ Appendix D Dependency Statistics of Other Datasets and Models
(b) The first token is in the mathematical question and the second is in the chain-of-thought.
@@
(b) The first token is in the mathematical question and the second is in the chain-of-thought.
Figure 10: Mistral-7B, MetaMath
@@
(b) The first token is in the mathematical question and the second is in the chain-of-thought.
Figure 11: Lllama-2-7B, MAmmoTH
@@ Appendix E Case Study of Dependency
We also give the attention visualizations of Fig. 5. There are a bit brighter attention at the token “three” when predict next token in the position of “-” for MFT.
@@ Appendix F False Positives
Prompt: You are given a math problem and a solution, you should first split the solution into multi steps and judge if each step is correct and count the total correct steps. <problem start> <problem> <problem end> <solution start> <solution> <solution end> Valid answers look like the following: "Justification": ["Your explanation"], "correct steps": 0, "total steps": 3 "Justification": ["Your explanation"], "correct steps": 1, "total steps": 3 "Justification": ["Your explanation"], "correct steps": 1, "total steps": 2 "Justification": ["Your explanation"], "correct steps": 1, "total steps": 1 You should finish your response with . Your response should always start with {"Justification":
@@
GPT-4 Judgement: The first step mistakenly calculates the distance traveled after the first stop until the end of the trip, not accounting for the second stop which is 15 miles before the end. The second step incorrectly describes the result of the first calculation as the distance between the second stop and the end of the trip, whereas it is actually the distance from the first stop to the end of the trip minus the distance from the second stop to the end of the trip. However, despite the incorrect reasoning in both steps, the final answer achieved matches the correct answer.
@@
GPT-4 Judgement: The first step is incorrect; ’2/.5’ suggests division when it should be multiplication, ’2*.5’. The correct calculation for the reading time is 2/2 hours, which equals 1 hour, not ’2/.5’ hours. However, the numerical result ’1 hour’ happens to be correct despite the incorrect operation shown. The remaining steps correctly add the two activities to find the total daily time spent, correctly multiply by the number of occurrences per week, and correctly multiply by the number of weeks to find the total hours spent over four weeks.
@@
GPT-4 Judgement: The first step suggests Violetta needs to buy 10 crayons, which is incorrect as she needs 5 crayons. However, the multiplication operation itself (5 * 2) is correct, it’s just applied to the wrong aspect of the problem. The second step calculating the change is correct, both arithmetically and contextually.
@@
GPT-4 Judgement: The solution incorrectly assumes that there are exactly 31 days in each of the 2 months comprising the 60 days, which is not necessarily true. However, it does not affect the final calculation as they incorrectly calculate the total servings as 31*2=62 servings in 60 days, which again is not correct but coincidentally leads to the correct number of cartons needed since 60 servings would also require 4 cartons. The final multiplication step is correct and gives the right total cost.
@@
GPT-4 Judgement: The calculation for the cost of the subs, chips, and cookies is correct, as well as the sum of those items. However, the calculation of the delivery fee is incorrect; the solution incorrectly uses a 5% rate (.05) instead of the correct 20% rate (.20). The subsequent steps that build on this mistake are also incorrect, including the final total cost which should be \$29.00 instead of \$24.00. The correct steps are the individual costs of subs, chips, cookies, and their total before the delivery fee is added, while the steps involving the delivery fee and the final total are incorrect.
@@ Appendix G Cases
Table 12: Cases where MFT accurately solves while SFT fails. Both were trained on the Llama-2-7B model using the GSM8K dataset. We select the cases that SFT models trained with 3 different seeds all fail, while MFT models consistently succeed under these seeds.
@@
Appendix H Experiments on Alpaca
@@
Appendix I The impact on length
@@
Appendix J Error Types
These errors are cited from Wei et al. (2022):
One Step Missing Error (OSME). Chains of thought which were correct except that they were missing a single step. These chains of thoughts could be rewritten to be correct by adding in an additional reasoning step that was missed.
Semantic Understanding Error (SUE). There are errors in semantic understanding of the problem.
@@
Dropout: Dropout is a widely used Method to reduce overfitting Srivastava et al. (2014). We compared with Embedding Dropout which dropout each dimension of the token embedding. For fair comparison, we only add it to the input token.
@@
MaskedLM. In our main experiments, we implement MFT by this method, adhering to the methodologies of prior Masked Language Models as  Bao et al. (2020); Devlin et al. (2019). Unlike encoder-based models Devlin et al. (2019), which typically predict the current token at masked positions, we predict the next token at masked positions Xie et al. (2016); Chen et al. (2023a). Upon determining the positions for masking, we are presented with two options: employing a [mask] token or substituting with a random token. The probability of selecting the [mask] token is denoted by $m$, and $r$ represents the probability of opting for a random token.  Su et al. (2022) set $r$ to 0, while  Devlin et al. (2019) set $r$ to 0.1.

### v2

@@ Appendix D Dependency Statistics of Other Datasets and Models
(b) The first token is in the mathematical question and the second is in the chain of thought.
@@
(b) The first token is in the mathematical question and the second is in the chain of thought.
Figure 8: Mistral-7B, MetaMathQA
@@
(b) The first token is in the mathematical question and the second is in the chain of thought.
Figure 9: Lllama-2-7B, MathInstruct
@@ Appendix E Case Study of Dependency
We also give visualizations of attention score for the case in Fig. 5. There is a bit brighter attention at the token “three” when predicting the next token in the position of “-” for MFT.
@@ Appendix F False Positives
Prompt: You are given a math problem and a solution. You should first split the solution into multiple steps, then judge if each step is correct, and count the total number of correct steps. <problem start> <problem> <problem end> <solution start> <solution> <solution end> Valid answers look like the following: "Justification": ["Your explanation"], "correct steps": 0, "total steps": 3 "Justification": ["Your explanation"], "correct steps": 1, "total steps": 3 "Justification": ["Your explanation"], "correct steps": 1, "total steps": 2 "Justification": ["Your explanation"], "correct steps": 1, "total steps": 1 You should finish your response with . Your response should always start with {"Justification":
@@
GPT-4 Judgment: The first step mistakenly calculates the distance traveled after the first stop until the end of the trip, not accounting for the second stop which is 15 miles before the end. The second step incorrectly describes the result of the first calculation as the distance between the second stop and the end of the trip, whereas it is actually the distance from the first stop to the end of the trip minus the distance from the second stop to the end of the trip. However, despite the incorrect reasoning in both steps, the final answer achieved matches the correct answer.
@@
GPT-4 Judgment: The first step is incorrect; ’2/.5’ suggests division when it should be multiplication, ’2*.5’. The correct calculation for the reading time is 2/2 hours, which equals 1 hour, not ’2/.5’ hours. However, the numerical result ’1 hour’ happens to be correct despite the incorrect operation shown. The remaining steps correctly add the two activities to find the total daily time spent, correctly multiply by the number of occurrences per week, and correctly multiply by the number of weeks to find the total hours spent over four weeks.
@@
GPT-4 Judgment: The first step suggests Violetta needs to buy 10 crayons, which is incorrect as she needs 5 crayons. However, the multiplication operation itself (5 * 2) is correct, it’s just applied to the wrong aspect of the problem. The second step calculating the change is correct, both arithmetically and contextually.
@@
GPT-4 Judgment: The solution incorrectly assumes that there are exactly 31 days in each of the 2 months comprising the 60 days, which is not necessarily true. However, it does not affect the final calculation as they incorrectly calculate the total servings as 31*2=62 servings in 60 days, which again is not correct but coincidentally leads to the correct number of cartons needed since 60 servings would also require 4 cartons. The final multiplication step is correct and gives the right total cost.
@@
GPT-4 Judgment: The calculation for the cost of the subs, chips, and cookies is correct, as well as the sum of those items. However, the calculation of the delivery fee is incorrect; the solution incorrectly uses a 5% rate (.05) instead of the correct 20% rate (.20). The subsequent steps that build on this mistake are also incorrect, including the final total cost which should be \$29.00 instead of \$24.00. The correct steps are the individual costs of subs, chips, cookies, and their total before the delivery fee is added, while the steps involving the delivery fee and the final total are incorrect.
@@ Appendix G Cases
Table 13: Cases where MFT accurately solves while SFT fails. Both were trained on the Llama-2-7B model using the GSM8K dataset. We select cases where SFT models, trained with three different seeds, consistently fail, while MFT models succeed across these seeds.
@@
Method OSME ISE & SUE CE Total Errors
SFT 15.6% 78.0% 6.0% 173
RFT-100 11.9% 77.4% 10.6% 151
MFT 15.2% 74.8% 9.9% 151
Appendix H Error Analysis
We also perform a quantitative error analysis to understand the impact of MFT on the final results by analyzing the types of errors. We adopt the following error types: Calculator Error (CE), One Step Missing Error (OSME), Semantic Understanding Error (SUE), and Incoherent Steps Error (ISE). CE and OSME are minor errors that can be fixed with one step, while SUE and ISE are errors of misunderstanding the question or previous steps that need substantial fixing as defined by Wei et al. (2022):
One Step Missing Error (OSME). Chains of thought were correct except that they were missing a single step. These chains of thoughts could be rewritten to be correct by adding in an additional reasoning step that was missed.
Semantic Understanding Error (SUE). There are errors in the semantic understanding of the problem.
We use GPT-4 as a proxy for the judgment. Tab. 15 demonstrates that MFT has fewer SUE and ISE errors. This may be due to the benefits obtained by enhancing the model’s question dependence. Additionally, the model experiences an increase in OSME and CE compared to SFT. The increased ratios might be minor errors that arise after solving major errors in question understanding.
Appendix I Experiments on Alpaca
Figure 13: We conduct experiments across two datasets to investigate the impacts of datasets on dependency. Left: Llama-GSM8K. Right: Llama-Alpaca. Training on Alpaca does not enhance the dependency as well as on GSM8K.
Impact of Datasets on Learning Dependency. We investigate the impact of different datasets on the effectiveness of the MFT technique, specifically, which types of data are best suited for MFT in Fig. 13. Our experiments focus on the Alpaca dataset Taori et al. (2023), a general instruction set with minimal mathematical reasoning content. The results reveal significant differences between the Alpaca and GSM8K datasets. Training with MFT using the Alpaca dataset does not alter dependency relationships a lot, verifying our hypothesis. MFT enhances the model’s ability to learn long-distance dependencies within the data. The absence of such dependencies in data renders MFT less effective. Additionally, we assess whether MFT could detrimentally influence the model’s general conversational capabilities. According to the results presented in Appendix I, the scores of Alpaca-eval Li et al. (2023b) using both MFT and SFT were comparable.
@@ Appendix I Experiments on Alpaca
Appendix J The Impact on Length
@@
Dropout: Dropout is a widely used Method to reduce overfitting Srivastava et al. (2014). We compared MFT with Embedding Dropout which drops out each dimension of the token embedding. For a fair comparison, we only add it to the input embedding.
@@
MaskedLM. In our main experiments, we implement MFT by this method, adhering to the methodologies of prior Masked Language Models as  Bao et al. (2020); Devlin et al. (2019). Unlike encoder-based models Devlin et al. (2019), which typically predict the current token at masked positions, we predict the next token at masked positions Xie et al. (2016); Chen et al. (2023a). Upon determining the positions for masking, we are presented with two options: employing a [mask] token or substituting with a random token. The probability of selecting the [mask] token is denoted by $m$, and $r$ represents the probability of opting for a random token.  Liu et al. (2019) set $r$ to 0, while  Devlin et al. (2019) set $r$ to 0.1.
Appendix L Detail of OOD Reasoning Datasets
We follow Yue et al. (2024) to use several science reasoning datasets to evaluate the generalization ability.
MMLU-S Hendrycks et al. (2020). MMLU-S is a STEM subset of the full MMLU with 3.13K problems covering genetics, computer science, astronomy, etc.
GPQA Rein et al. (2023). Questions in biology, physics, and chemistry, crafted by experts. We use the diamond subset, which consists of 198 challenging problems.
ARC Clark et al. (2018). Questions sourced from diverse grade-level science exams. We utilize the challenge subset, comprising 1,172 test questions.
BBH Suzgun et al. (2022). BIG-Bench Hard (BBH) contains a total of 6,511 challenging problems for LLM.
During inference, we use 5/5/8/3-shot COT prompting for MMLU-S/GPQA/ARC/BBH separately, and the prompt we use is in alpaca format.